OpenStudy (anonymous):

the ratio of frogs to fish in a pond is 4:21. there r 84 fish in the pond. how many frogs r in the pond?

OpenStudy (anonymous):

\[\frac{4}{21}=\frac{x}{84}\iff x=\frac{84\times 4}{21}\]

OpenStudy (anonymous):

what

OpenStudy (anonymous):

ratio is the same as a fraction \[\text {ratio of frogs to fish}=4:21\] is the same as saying \[\frac{\text{# of frogs}}{\text{# of fish}}=\frac{4}{21}\]

OpenStudy (anonymous):

you know the number of fish is 84 you don't know the number of frogs set up the ratio as \[\frac{4}{21}=\frac{\text{ # of frogs}}{84}\]

OpenStudy (anonymous):

solve this for the number of frogs by computing \[\text{# of frogs}=\frac{84\times 4}{21}\]

OpenStudy (anonymous):

thanks